# Learning a Neural Solver for Parametric PDE to Enhance Physics-Informed Methods

### v2

@@ E.2 Statistical analysis
We provide a statistical analysis on some datasets. These experiments are conducted using $L=5$ (instead of $L=2$ in table 2). For computational cost reasons, we did not make this experiment on all datasets, but kept $2$ datasets in $1d$ (table 11), 1 dataset in $1$d+time and 1 dataset in $2$d (table 12) so that several configurations and data sizes are represented.
@@
Table 11: Results of trained models with error bars on 1d datasets - metrics in Relative MSE on test set. Best performances are highlighted in bold, second best are underlined.
@@
Table 12: Results of trained models with error bars on 1d + time and 2d datasets - metrics in Relative MSE on test set. Best performances are highlighted in bold, second best are underlined.
@@ E.3 Computational cost
As we can see on table 13, our model takes longer to train due to the iterative process occurring at each epoch. However, note that this is training time; inference time is similar to other methods (see table 14). We detail a justification for this additional training time compared to PINNs variants below:
  • Comparison with vanilla PINNs: Consider the following experiment. Suppose we train a classical PINN on a single instance of the Darcy PDE. Based on the training times shown in table 13, using this method, we performed $1,500$ steps, which took $420$ minutes for training (please note that the performances were less accurate than our model’s performance). If we wanted to train a PINNs on each PDE of out entire test dataset for $15000$ epochs (sometimes even more steps are required), this would take $4200$ minutes or stated otherwise approximatly $3$ days. Suppose now, that one wants to solve an additional equation. This will require an average of $0.226$ seconds (see table 14) with our method, while PINNs would require an entirely new training session of approximately $20$ minutes for $15,000$ steps. This makes our method $5,000$ times faster than traditional PINNs for solving any new equation.
  • Comparison with PINNs parametric variants: Now let us consider a parametric variant of PINNs to handle multiple PDEs (PPINNs for parametric PINNs). In table 2, PPINNs was trained for $5000$ epochs only. Let us consider training it further as suggested for vanilla PINNs, until $15,000$ epochs. First, this training would require approximately $19$h30303030m. Then, the optimization problem is still ill-conditioned, training further would probably not significantly improve the performance.
@@
Table 13: Training time of the experiments shown in Table 2. of the paper on a NVIDIA TITAN RTX (25 Go) GPU. d stands for days, h for hours, m for minutes.
@@
Table 14: Inference time on the test set (in seconds). All experiment are conducted on a single NVIDIA RTX A6000 (48Go). We report the mean time computed on the test set to evaluate the baselines as performed on Table 2 in the paper (i.e. with $10$ test-time optimization steps when applicable and $20$ steps on the Heat dataset).
@@ E.4 Additional datasets
We provide additional experiment on $2$ new datasets: Non-linear Reaction-Diffusion in a more complex setting and Advections. We refer the reader to appendix D for the details about the PDE setting. These datasets were not included in the main part of the paper due to a lack of space in the results table.
@@
Table 15: Results of trained models on additional datasets - metrics in Relative MSE on test set. Best performances are highlighted in bold, second best are underlined.
@@
These additional dataset show cases where baselines are performing very well on the considered PDE. For Advections, PINO reached very good performances. We believe that the Fourier layers used in the model fits well to the phenomenon. Indeed, the solution is represented using a combination of sine moving with time. This simple dynamics is easily captured using Fourier transformations. Our B-Splines basis can be sub-optimal for this dataset. For the complexified version of NLRD, NLRDIC, it is the supervised baseline that performs best. This dataset, does not present high frequencies, and the MLP looks sufficiently expressive to find the coefficient in the basis. However, this baseline had difficulties to reconstruct the higher frequencies in Advections. Even if our model does not perform best on these datasets, it is ranked second. We believe that these additional results show the robustness of our method.
@@ Appendix F Qualitative results
This section is dedicated to visualization of the results of our model, baselines and optimizers, presented in section 4. For each PDE considered, we chose $2$ samples in the test sets and compute the solutions with our model and the different baselines. We provide visualization samples with $L=5$ i.e. results proposed in tables 11 and 12 to detail more precisely the evolution of the solution at several steps of optimization. $3$ datasets are shown with $L=2$: Heat for computational reasons (training with $L=5$ is much more expensive when the dimension of the problem increases) and the $2$ additional datasets trained only with $L=2$ (Advections and NLRDIC). Then, we show the evolution of the reconstruction of the solution with our method i.e. we plot the solution at each step of the optimization (figs. 13(a), 13(b), 16(a), 16(b), 19(a), 19(b), 22(a), 22(b), 25(a), 25(b), 28(a), 28(b), 31(a) and 31(b)) and we compare the final prediction with baselines’ (figs. 14(a), 14(b), 17(a), 17(b), 20(a), 20(b), 23(a), 23(b), 26(a), 26(b), 29(a), 29(b), 32(a) and 32(b)). Finally, we chose $20$ ($6$ for Heat) PDEs and we reproduce fig. 2 for every dataset. More precisely, we optimize one PINN per PDE using Adam, we fit our basis using several optimizers (GD, ADAM, L-BFGS and our learned optimization process) and we fine-tune the learned PINO for $10,000$ steps and visualize the evolution of the MSE (averaged at each step on the selected PDE). These figures show the relevance of learning the optimizer when using physics-informed losses (figs. 15, 18, 21, 24, 27, 30 and 33).

### v3

@@ E.2 Statistical analysis
We provide a statistical analysis on some datasets and baselines. These experiments are conducted using $L=5$ (instead of $L=2$ in table 2). For computational cost reasons, we did not make this experiment on all datasets but kept $2$ datasets in $1d$ (table 14), 1 dataset in $1$d+time, and 1 dataset in $2$d (table 15) so that several configurations and data sizes are represented.
@@
Table 14: Results of trained models with error bars (std errors) on 1d datasets - metrics in Relative MSE on test set. Best performances are highlighted in bold, and second best are underlined.
@@
Table 15: Results of trained models with error bars (std errors) on 1d + time, and 2d datasets - metrics in Relative MSE on test set. Best performances are highlighted in bold, and second best are underlined.
This analysis shows the robustness of our proposed method w.r.t. initial seed.
@@ E.3 Computational cost
Finally, we detail here the training and inference times of our method as well as baselines.
As we can see in table 16, our model takes longer to train due to the iterative process occurring at each epoch. However, note that this is training time; inference time is similar to other methods (see table 17). We detail a justification for this additional training time compared to PINNs variants below:
  • Comparison with vanilla PINNs: Consider the following experiment. Suppose we train a classical PINN on a single instance of the Darcy PDE. Based on the training times shown in table 16, using this method, we performed $1,500$ steps, which took $420$ minutes for training (please note that the performances were less accurate than our model’s performance). If we wanted to train a PINNs on each PDE of our entire test dataset for $15000$ epochs (sometimes even more steps are required), this would take $4200$ minutes or stated otherwise approximatly $3$ days. Suppose now, that one wants to solve an additional equation. This will require an average of $0.226$ seconds (see table 17) with our method, while PINNs would require an entirely new training session of approximately $20$ minutes for $15,000$ steps. This makes our method $5,000$ times faster than traditional PINNs for solving any new equation.
  • Comparison with PINNs parametric variants: Now let us consider two parametric variants of PINNs designed to handle multiple PDEs (PPINNs for parametric PINNs, P2INNs for the model proposed by (Cho et al., 2024)). In table 2, models was trained for $5,000$ epochs only. Let us consider training it further as suggested for vanilla PINNs, until $15,000$ epochs. First, this training would require approximately $19$h30303030m. Then, the optimization problem is still ill-conditioned, training further would probably not significantly improve the performance. We can extend this reasoning to the P2INNs baseline, for which we observed similar performance and behaviors.
@@
Table 16: Training time of the experiments shown in Table 2. of the paper on a single NVIDIA TITAN RTX (25 Go) GPU. d stands for days, h for hours, m for minutes.
@@
Table 17: Inference time, averaged (in seconds) on the test set. All experiments are conducted on a single NVIDIA RTX A6000 (48Go). We report the mean time computed on the test set to evaluate the baselines as performed in table 2 in the paper (i.e. with $10$ test-time optimization steps when applicable and $20$ steps on the Heat dataset). We consider as inference the solving of a PDE given its parameters and/or initial/boundary conditions.
@@ E.4 Additional datasets
We provide additional experiments on $2$ new datasets: Non-linear Reaction-Diffusion in a more complex setting and Advections. We refer the reader to appendix C for the details about the PDE setting. These datasets were not included in the main part of the paper due to a lack of space in the results table.
@@
Table 18: Results of trained models on additional datasets - metrics in Relative MSE on the test set. Best performances are highlighted in bold, and second best are underlined.
@@
These additional datasets show cases where baselines are performing very well on the considered PDE. For Advections, PINO reached very good performances. We believe that the Fourier layers used in the model fit well to the phenomenon. Indeed, the solution is represented using a combination of sine moving with time. This simple dynamics is easily captured using Fourier transformations. Our B-Splines basis can be sub-optimal for this dataset. For the complexified version of NLRD, NLRDIC, it is the supervised baseline that performs best. This dataset does not present high frequencies, and the MLP looks sufficiently expressive to find the coefficient in the basis. However, this baseline had difficulties in reconstructing the higher frequencies in Advections. Even if our model does not perform best on these datasets, it is ranked second. We believe that these additional results show the robustness of our method across different physical phenomena.
E.5 Training behavior
We show in this section the evolution of the MSE as the training progresses (see fig. 15).
Figure 15: MSE during training on the training and testing sets. Example shown on the Helmholtz equation for results as presented in table 2.
Figure 16: Reconstruction of the solution PDEs during training. Example drawn from the test set.
In fig. 15, we show the MSE evolution on both the train and test sets (evaluation every $50$ epochs) and on figs. 16(a) and 16(b) some reconstruction (from the test set), with respect to the training epochs. The training set ($800$ PDE trajectories) corresponds to PDEs used to update the network parameters, while the test set ($200$ trajectories) are unseen PDEs for the network. This means that the model has not been trained or optimized on these PDEs. The test set is composed of PDEs with varying PDE parameter values ($\omega$) and boundary conditions ($u_{0},v_{0}$). This illustrates that the generalization performance on new PDEs within the training distribution is rapidly achieved by the network.
E.6 Loss landscapes
In this section, we propose a visual representation of the optimization paths in the loss landscape. Figures 17 and 18 illustrates the behavior of the vanilla PINNs algorithm and of our "learning to optimize" method. The plots represent respectively a 2D visualization of the physical loss landscape $\mathcal{L}_{PDE}$ (fig. 17) and of the data loss landscape $\mathcal{L}_{\text{DATA}}$ (fig. 18), around the approximate solution (in our basis) of a given Helmholtz equation. The function basis used in this experiment has a size of $32$. We used the technique in (Garipov et al., 2018) to visualize the solution coordinates in the basis $\Theta$ in the loss landscape (more details in section E.7.1). Superimposed on the loss background, we plot three trajectories obtained by starting from an initial random vector $\Theta_{0}$.
Figure 17 visualizes the sharp and ill conditioned landscape of the physical loss $\mathcal{L}_{PDE}$, while Figure 18 shows the better conditioned landscape of the MSE data loss $\mathcal{L}_{\text{DATA}}$. The figures provide intuition on how the proposed algorithm operates and improves the convergence. The optimizer $\mathcal{F}$ modifies the direction and magnitude of the stochastic gradient descent (SGD) updates in the physical loss landscape (Figure 17, left plot). It achieves this by utilizing solution values at various sample points to adjust the gradient, steering it toward the corresponding minimum in the mean squared error (MSE) landscape (Figure 18, central plot). While the descent remains within the residual physics error space, the learned optimizer provides an improved gradient direction, enhancing convergence efficiency (Figure 17, right plot). We describe below the different figures.
  • Left Columns in Figures 17 and 18: these figures show the gradient path ($100$ steps) obtained by directly optimizing the physical loss $\mathcal{L}_{PDE}$, similar to the vanilla PINNs algorithm. This trajectory highlights the ill-conditioning of the optimization problem associated to $\mathcal{L}_{PDE}$.
  • Center Column in Figures 17 and 18: these figures plot the trajectory of a gradient-based optimization algorithm trained with a mean squared error $\mathcal{L}_{\text{DATA}}$ data loss ($100$ steps), under the assumption that the solution values are known at collocation points. While these quantities are not available in our case, this visualization is included to illustrate the differences in convergence behavior between physics-based and MSE-based loss functions when both are accessible.
  • Right Column in Figures 17 and 18: these plots illustrate the behavior of our algorithm (the solver is trained with 2 gradient steps for this example). It demonstrates the effect of our learned optimizer and the significant improvements achieved compared to a standard gradient descent algorithm on the physical loss.
. We describe in section E.7 the construction of the figures figs. 17 and 18.
Figure 17: Loss landscapes and optimization trajectories on an instance of the Helmholtz PDE. The figure represents the PDE loss $\mathcal{L}_{\text{PDE}}$ landscape and superimposed are examples of optimization paths computed using SGD on the physical loss (left), the data loss (center), and our method (right).
Figure 18: Loss landscapes and optimization trajectories on an instance of the Helmholtz PDE. The figures represents the DATA loss $\mathcal{L}_{\text{DATA}}$ and superimposed are examples of optimization paths computed using SGD on the physical loss (left), the data loss (center), and our method (right).
E.7 Projection of the basis for visualization
Following the technique described in (Garipov et al., 2018), we plot a 2D slice of the loss function around the minimum solution. We explain below the method used for creating figs. 17 and 18.
E.7.1 Create a basis for projection
We want to visualize the loss landscapes around a parameter solution $\Theta^{\star}$ of a given PDE. Due to the high dimensionality of the parameter space, we use the method proposed in (Garipov et al., 2018), that involves creating a 2d slice of the loss landscape for visualization. This methods projects the multidimensional landscape onto a 2D basis $(u,v)$ so that any solution vector $\Theta$ could be written as $\Theta=\alpha u+\beta v$, with, $\alpha,\beta\in\mathbb{R}$ and $u,v\in\mathbb{R}^{N}$, $2$ orthonormal vectors. Let $\Omega_{u,v}:=\text{Vect}\{u,v\}$ denote this projection space. For the visualization, the basis must include the solution vectors found by our algorithm. For that:
  1. 1.
    We start by running our optimization algorithm from a random vector parameters $\Theta_{0}$ to find $\Theta_{L}\in\mathbb{R}^{N}$, our approximate PDE solution. We set $w_{1}=\Theta_{L}$ and set a second vector $w_{2}=\Theta_{0}$, chosen here as the initial vector of our optimization process.
  2. 2.
    We generate a third vector $w_{3}\in\mathbb{R}^{N}$ randomly.
  3. 3.
    We then set: $u=w_{2}-w_{1}$ and $v=(w_{3}-w_{1})-\frac{<w_{3}-w_{1},w_{2}-w_{1}>}{||w_{2}-w_{1}||^{2}}(w_{2}-w_{1})$
  4. 4.
    The resulting normalized vectors $\hat{u}=\frac{u}{||u||}$ and $\hat{v}=\frac{v}{||v||}$ form an orthogonal basis containing $w_{1},w_{2}$ and $w_{3}$.
  5. 5.
    Using the solution parameters $\Theta_{L}$ as the origin for our newly created basis, we can compute a $2$D slice of the loss landscape where each solution $\Theta$ could be represented in the $(u,v)$ basis with its projection coordinates $(\alpha,\beta)$ as $\Theta_{projection}=w_{1}+\alpha\hat{u}+\beta\hat{v}$.
E.7.2 Plot trajectories
Finally, we run a baseline optimization procedures and get trajectories $\{\Theta_{i}\}_{i=0}^{L}$, with L being the number of optimization steps, e.g. a gradient descent, on the physical loss, the MSE loss or our proposed method. We project the $\Theta_{i}$’s in the created basis (see section E.7.1) and plot this projection of the optimization trajectory on the loss landscape $\Omega_{u,v}$ as visualized on figs. 17 and 18.
E.7.3 Illustration of the ill-conditioning of the PDE loss
As an illustration of the ill-conditioning of the PDE loss, we replicate fig. 17 using a basis that further emphasizes this aspect. To build this basis, we use the procedure described in section E.7.1 by using $2$ eigenvectors of the hessian of the PDE loss. First, we compute $\text{Hess}(\mathcal{L}_{\text{PDE}})$ and its eigenvectors decomposition. We select the vectors associated to the highest and lowest eigenvalues, and respectively set them to $w_{2}$ and $w_{3}$. The resulting landscape visualization is shown in fig. 19.
Figure 19: Loss landscapes and optimization trajectories on an instance of the Helmholtz PDE. These optimization trajectories are computed using SGD on the physical loss (left), the data loss (center), and our method (right). The background represents the PDE loss $\mathcal{L}_{\text{PDE}}$.
Figure 19, clearly illustrates the characteristic shape of the ill-conditioned function $\mathcal{L}_{\text{PDE}}$. The two directions, extracted from the highest and lowest eigenvalues, are clearly visible on this PDE loss landscape. This highlights the difficulty of this optimization problem for standard descent methods.
@@ Appendix F Qualitative results
This section is dedicated to visualization of the results of our model, baselines and optimizers, presented in section 4. For each PDE considered, we chose $2$ samples in the test sets and compute the solutions with our model and the different baselines. We provide visualization samples with $L=5$ i.e. results proposed in tables 14 and 15 to detail more precisely the evolution of the solution at several steps of optimization. $3$ datasets are shown with $L=2$: Heat for computational reasons (training with $L=5$ is much more expensive when the dimension of the problem increases) and the $2$ additional datasets trained only with $L=2$ (Advections and NLRDIC). Then, we show the evolution of the reconstruction of the solution with our method i.e. we plot the solution at each step of the optimization (figs. 20(a), 20(b), 23(a), 23(b), 26(a), 26(b), 29(a), 29(b), 32(a), 32(b), 35(a), 35(b), 38(a) and 38(b)) and we compare the final prediction with baselines’ (figs. 21(a), 21(b), 24(a), 24(b), 27(a), 27(b), 30(a), 30(b), 33(a), 33(b), 36(a), 36(b), 39(a) and 39(b)). Finally, we chose $20$ ($6$ for Heat) PDEs and we reproduce fig. 2 for every dataset. More precisely, we optimize one PINN per PDE using Adam, we fit our basis using several optimizers (GD, ADAM, L-BFGS and our learned optimization process) and we fine-tune the learned PINO for $10,000$ steps and visualize the evolution of the MSE (averaged at each step on the selected PDE). These figures show the relevance of learning the optimizer when using physics-informed losses (figs. 22, 25, 28, 31, 34, 37 and 40).